# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2024-125 LHCb-PAPER-2024-007

### v1

@@ H. Wu67
Z. Xiang5
Y. Xie8
A. Xu33
J. Xu7
L. Xu4
L. Xu4
M. Xu55
Z. Xu11
Z. Xu7
Z. Xu5
D. Yang4
K. Yang60
S. Yang7
X. Yang6
Y. Yang27,n
Z. Yang6
Z. Yang65
V. Yeroshenko13
H. Yeung61
H. Yin8
C. Y. Yu6
J. Yu69
X. Yuan5
E. Zaffaroni48
M. Zavertyaev19
M. Zdybal39
C. Zeng5,7
M. Zeng4
C. Zhang6
D. Zhang8
J. Zhang7
L. Zhang4
S. Zhang69
S. Zhang6
Y. Zhang6
Y. Z. Zhang4
Y. Zhao20
A. Zharkova42
A. Zhelezov20
X. Z. Zheng4
Y. Zheng7
T. Zhou6
X. Zhou8
Y. Zhou7
V. Zhovkovska55
L. Z. Zhu7
X. Zhu4
X. Zhu8
V. Zhukov16
J. Zhuo46
Q. Zou5,7
D. Zuliani31,q
G. Zunica48
@@
2Centro Brasileiro de Pesquisas Físicas (CBPF), Rio de Janeiro, Brazil
@@
10Université Savoie Mont Blanc, CNRS, IN2P3-LAPP, Annecy, France
11Université Clermont Auvergne, CNRS/IN2P3, LPC, Clermont-Ferrand, France
@@
13Université Paris-Saclay, CNRS/IN2P3, IJCLab, Orsay, France
@@
15LPNHE, Sorbonne Université, Paris Diderot Sorbonne Paris Cité, CNRS/IN2P3, Paris, France
@@
17Universität Bonn - Helmholtz-Institut für Strahlen und Kernphysik, Bonn, Germany
@@
19Max-Planck-Institut für Kernphysik (MPIK), Heidelberg, Germany
@@
38AGH - University of Krakow, Faculty of Physics and Applied Computer Science, Kraków, Poland
39Henryk Niewodniczanski Institute of Nuclear Physics Polish Academy of Sciences, Kraków, Poland
@@
45Instituto Galego de Física de Altas Enerxías (IGFAE), Universidade de Santiago de Compostela, Santiago de Compostela, Spain
@@
48Institute of Physics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Lausanne, Switzerland
49Physik-Institut, Universität Zürich, Zürich, Switzerland
@@
68Pontifícia Universidade Católica do Rio de Janeiro (PUC-Rio), Rio de Janeiro, Brazil, associated to 3
@@
78Universidade da Coruña, A Coruna, Spain, associated to 43
@@
aUniversidade de Brasília, Brasília, Brazil
bCentro Federal de Educacão Tecnológica Celso Suckow da Fonseca, Rio De Janeiro, Brazil
@@
eLIP6, Sorbonne Université, Paris, France
@@
gUniversidad Nacional Autónoma de Honduras, Tegucigalpa, Honduras
hUniversità di Bari, Bari, Italy
@@
jUniversità di Bologna, Bologna, Italy
kUniversità di Cagliari, Cagliari, Italy
lUniversità di Ferrara, Ferrara, Italy
mUniversità di Firenze, Firenze, Italy
nUniversità di Genova, Genova, Italy
oUniversità degli Studi di Milano, Milano, Italy
pUniversità degli Studi di Milano-Bicocca, Milano, Italy
qUniversità di Padova, Padova, Italy
rUniversità di Perugia, Perugia, Italy
@@
tUniversità di Pisa, Pisa, Italy
uUniversità della Basilicata, Potenza, Italy
vUniversità di Roma Tor Vergata, Roma, Italy
wUniversità di Siena, Siena, Italy
xUniversità di Urbino, Urbino, Italy
yUniversidad de Alcalá, Alcalá de Henares , Spain

### v2

@@ H. Wu67â
Z. Xiang5â
Y. Xie8â
A. Xu33â
J. Xu7â
L. Xu4â
L. Xu4â
M. Xu55â
Z. Xu11â
Z. Xu7â
Z. Xu5â
D. Yang4â
K. Â Yang60â
S. Yang7â
X. Yang6â
Y. Yang27,nâ
Z. Yang6â
Z. Yang65â
V. Yeroshenko13â
H. Yeung61â
H. Yin8â
C. Y. Â Yu6â
J. Yu69â
X. Yuan5â
E. Zaffaroni48â
M. Zavertyaev19â
M. Zdybal39â
C. Â Zeng5,7â
M. Zeng4â
C. Zhang6â
D. Zhang8â
J. Zhang7â
L. Zhang4â
S. Zhang69â
S. Zhang6â
Y. Zhang6â
Y. Z. Â Zhang4â
Y. Zhao20â
A. Zharkova42â
A. Zhelezov20â
X. Z. Â Zheng4â
Y. Zheng7â
T. Zhou6â
X. Zhou8â
Y. Zhou7â
V. Zhovkovska55â
L. Z. Â Zhu7â
X. Zhu4â
X. Zhu8â
V. Zhukov16â
J. Zhuo46â
Q. Zou5,7â
D. Zuliani31,qâ
G. Zunica48â
@@
2Centro Brasileiro de Pesquisas FĂsicas (CBPF), Rio de Janeiro, Brazil
@@
10UniversitĂŠ Savoie Mont Blanc, CNRS, IN2P3-LAPP, Annecy, France
11UniversitĂŠ Clermont Auvergne, CNRS/IN2P3, LPC, Clermont-Ferrand, France
@@
13UniversitĂŠ Paris-Saclay, CNRS/IN2P3, IJCLab, Orsay, France
@@
15LPNHE, Sorbonne UniversitĂŠ, Paris Diderot Sorbonne Paris CitĂŠ, CNRS/IN2P3, Paris, France
@@
17Universität Bonn - Helmholtz-Institut fßr Strahlen und Kernphysik, Bonn, Germany
@@
19Max-Planck-Institut fĂźr Kernphysik (MPIK), Heidelberg, Germany
@@
38AGH - University of Krakow, Faculty of Physics and Applied Computer Science, KrakĂłw, Poland
39Henryk Niewodniczanski Institute of Nuclear Physics Polish Academy of Sciences, KrakĂłw, Poland
@@
45Instituto Galego de FĂsica de Altas EnerxĂas (IGFAE), Universidade de Santiago de Compostela, Santiago de Compostela, Spain
@@
48Institute of Physics, Ecole Polytechnique FĂŠdĂŠrale de Lausanne (EPFL), Lausanne, Switzerland
49Physik-Institut, Universität Zßrich, Zßrich, Switzerland
@@
68PontifĂcia Universidade CatĂłlica do Rio de Janeiro (PUC-Rio), Rio de Janeiro, Brazil, associated to 3
@@
78Universidade da CoruĂąa, A Coruna, Spain, associated to 43
@@
aUniversidade de BrasĂlia, BrasĂlia, Brazil
bCentro Federal de EducacĂŁo TecnolĂłgica Celso Suckow da Fonseca, Rio De Janeiro, Brazil
@@
eLIP6, Sorbonne UniversitĂŠ, Paris, France
@@
gUniversidad Nacional AutĂłnoma de Honduras, Tegucigalpa, Honduras
hUniversitĂ di Bari, Bari, Italy
@@
jUniversitĂ di Bologna, Bologna, Italy
kUniversitĂ di Cagliari, Cagliari, Italy
lUniversitĂ di Ferrara, Ferrara, Italy
mUniversitĂ di Firenze, Firenze, Italy
nUniversitĂ di Genova, Genova, Italy
oUniversitĂ degli Studi di Milano, Milano, Italy
pUniversitĂ degli Studi di Milano-Bicocca, Milano, Italy
qUniversitĂ di Padova, Padova, Italy
rUniversitĂ di Perugia, Perugia, Italy
@@
tUniversitĂ di Pisa, Pisa, Italy
uUniversitĂ della Basilicata, Potenza, Italy
vUniversitĂ di Roma Tor Vergata, Roma, Italy
wUniversitĂ di Siena, Siena, Italy
xUniversitĂ di Urbino, Urbino, Italy
yUniversidad de AlcalĂĄ, AlcalĂĄ de Henares , Spain